Please Help.
Use elimination to solve each system of equations.
3M-N=26 -2M-N=-24

Answer by checkley77(12844)

3M-N=26 MULTIPLY BY -1 & ADD.
-2M-N=-24
-3M+N=-26
------------------
-5M=-50
M=-50/-5
M=10 ANS.
3*10-N=26
30-N=26
-N=26-30
-N=-4
N=4 ANS.
PROOF:
-2*10-4=-24
-20-4=-24
-24=-24